# Jingyao Liu♠♣  Chen Huang♠♢†  Zhizhao GuanWenqiang Lei♠♣Yang Deng

### v1

@@ 5 Conclusion & Discussion
Unlike function-level tasks, project-level coding tasks offer a more comprehensive assessment of LLMs’ code generation capabilities. However, the inherent complexity of constructing robust project-level benchmarks presents significant challenges. A primary hurdle arises from the diverse and often unique code implementations LLMs produce for identical user requirements, which substantially increases the difficulty and cost associated with developing reliable automated evaluation mechanisms. Consequently, current project-level benchmarks, including the one presented in this paper, often feature limited dataset sizes (as detailed in Table 1).
Despite these challenges, our current work represents a crucial and reasonable starting point toward more rigorous project-level task evaluation. Our E2EDev marks a pivotal advancement in standardizing the benchmark of E2ESD frameworks by providing detailed requirement specifications and reliable evaluation protocols. While LLMs have demonstrated promise in code generation and function-level software development, our work highlights the significant challenges remaining in achieving fully automated E2ESD. Looking ahead, future work should focus more on designing effective, reliable, and cost-efficient E2ESD solutions that can truly realize the potential of LLMs in automating software development. Additionally, acknowledging the substantial resource investment required for such benchmark development, we are committed to continuously expanding our dataset and establishing a public leaderboard. This sustained effort is dedicated to fostering ongoing advancements in the application of LLMs for software development, pushing the boundaries of what is currently achievable.
6 Ethics statement
@@
Our work focuses solely on constructing the E2EDev benchmark for end-to-end software development tasks to evaluate the capabilities of LLM-based E2ESD frameworks. The study does not involve human subjects, sensitive personal data, or other ethical risks, and all source data are obtained from open-source communities. Any human annotation or verification of data was performed under controlled conditions; details are provided in the Appendix A. All experiments and analyses adhere to the ICLR Code of Ethics, and no procedures, datasets, or methods in this study raise concerns related to privacy, bias, or safety.
7 Reproducibility Statement
@@
To facilitate reproducibility, we publicly release the full E2EDev benchmark source code on GitHub (https://github.com/SCUNLP/E2EDev), along with the complete benchmark dataset on Hugging Face (https://huggingface.co/datasets/GuanZhiZhao/E2EDev). Detailed setup and usage instructions are provided in both Appendix D.3 and the repository README. All experiments reported in this paper—including dataset processing, model evaluation, and metric computation—can be fully reproduced using these resources.
Acknowledgements
@@
This work was supported in part by the National Natural Science Foundation of China (No. 62272330 and No.U24A20328); in part by the Fundamental Research Funds for the Central Universities (No. YJ202219); in part by the Science Fund for Creative Research Groups of Sichuan Province Natural Science Foundation (No. 2024NSFTD0035); in part by the National Major Scientific Instruments and Equipments Development Project of Natural Science Foundation of China under Grant (No. 62427820); in part by the Natural Science Foundation of Sichuan (No. 2024YFHZ0233); in part by the Singapore Ministry of Education (MOE) Academic Research Fund (AcRF) Tier 1 grant (No. MSS24C004).
@@ A.2.1 HITL-MAA Annotation Process
1Input: Source code annotated with test-ids;
2 Output: Executable and validated end-to-end test scripts with aligned requirements;
3
4/* Code Analyzer */;
@@ 6   Analyze frontend framework and interactive components;
8foreach JS file in project do
10
11Summarize the analysis for downstream modules;
12
13/* Requirement Extractor */;
@@ 14 Extract contextual requirements using both source code and analysis summary;
17foreach req in user_req_lst do
18   Human annotator validates and corrects the extracted requirement;
19
@@ 52         Human annotator intervenes to revise and finalize test script;
A.2.2 Protocols for Human-in-the-Loop & Manual Verification
Manual Verification Protocol & Inner-Agreement.
To ensure the consistency and quality of the free-text requirements and test case validation, we implement a post-hoc gold evaluation process. First, among five human annotators selected based on their expertise and professional experience in software testing, we designate one senior expert as the Gold Checker responsible for reviewing the annotations of the other annotators. This process employs an iterative annotation and feedback mechanism to resolve inconsistencies in a timely manner. Specifically, for both requirements and test case annotations, four regular annotators initially perform validation. For requirement annotations, annotators interact with the web application through the browser to experience and verify the functionality directly. For test cases, which are written in natural language and serve as step-by-step operational guides, annotators follow the instructions to verify the consistency and validity of the test cases against the requirements. If discrepancies or errors are detected, annotators revise the requirements or test case texts accordingly before submitting them to the Gold Checker for final judgment. The Gold Checker evaluates each annotation from three binary aspects: equivalence, completeness, and correctness, where a score of 1 indicates the criterion is met, and 0 otherwise. Only when the Fleiss’ Kappa score for all three aspects exceeds 0.7 does the Gold Checker consolidate a final version of the requirement or test case (which may include discarding it if necessary). If the agreement falls below the threshold, the Gold Checker provides detailed feedback highlighting the inconsistencies to the annotators, who then repeat the annotation process. If any annotator identifies a newly proposed test case, it must undergo the gold evaluation process. In such cases, all three validation aspects are initially scored as 0 by the annotators, and the Gold Checker reviews the new test case before returning it to all regular annotators for reconsideration and re-annotation. In cases where annotators do not recognize a newly proposed test case, they assign all three aspects a score of 0, and the Gold Checker returns the new test case for reconsideration and re-annotation by the regular annotators. This approach effectively balances thoroughness and efficiency, resulting in a reliable, scalable, and reproducible manual annotation protocol suitable for complex free-text requirement and test case validation tasks.
This rigorous validation workflow ensures high-quality, consistent annotations suitable for downstream tasks.
@@ A.2.3 Analysis on Human Labor during HITL-MAA
Table 5: Comparison of Manual vs. Semi-Automated Annotation
@@ A.3 Case Studies of E2EDev
E2ESD_Bench_06 """prompt""" You are tasked with implementing a complete web application using HTML, JavaScript, and CSS. Your implementation must strictly follow the specifications described below. ### SUMMARY # overview The Playable Piano web application allows users to interact with a virtual piano. Users can click on piano keys or press corresponding keyboard keys to play notes. The application includes features for adjusting volume and toggling the visibility of key labels. # predefined_options The web application initializes with a default volume level set to 0.5, and piano key labels are displayed by default upon launch. It features a virtual piano with both black and white keys, each mapped to specific keyboard characters. The black keys correspond to the uppercase letters W, E, T, Y, U, O, and P, while the white keys correspond to A, S, D, F, G, H, J, K, L, and ;. Each key element in the DOM is assigned a unique data-test-id attribute in the format data-test-id=\"piano-key-’x’\", where x is the lowercase version of the corresponding keyboard character (e.g., the key for W uses data-test-id=\"piano-key-’w’\"). # external_resources The corresponding audio files for each piano note are stored in the tunes directory, with filenames matching the lowercase key characters followed by .wav, such as tunes/a.wav, tunes/;.wav, tunes/d.wav, tunes/e.wav, and so on, covering all relevant keys including w, e, t, y, u, o, p, a, s, d, f, g, h, j, k, l, and ;. # external_js_libraries No external JavaScript libraries are used in this application. ### Functional Requirements Implement the following features as described. For each requirement, make sure the HTML structure, JavaScript behavior, and CSS styles match the specifications exactly. # REQUIREMENTS: - Requirement 1: When the user clicks any visible piano key on the webpage, identified by its data-testid (e.g., piano-key-a for the white key a or piano-key-w for the black key w’), the system must play the corresponding sound file from the tunes directory (e.g., tunes/a.wav for key a’). The clicked key must receive the active class to visually highlight it, and the highlight must be removed after 150ms by removing the active class. - Requirement 2:... - Requirement 3:... - Requirement 4:... - Requirement 5:...   """requirement_summary""" ### overview The Playable Piano web application allows users to interact with a virtual piano. Users can click on piano keys or press corresponding keyboard keys to play notes. The application includes features for adjusting volume and toggling the visibility of key labels. ### predefined_options The web application initializes with a default volume level set to 0.5, and piano key labels are displayed by default upon launch. It features a virtual piano with both black and white keys, each mapped to specific keyboard characters. The black keys correspond to the uppercase letters W, E, T, Y, U, O, and P, while the white keys correspond to A, S, D, F, G, H, J, K, L, and ;. Each key element in the DOM is assigned a unique data-test-id attribute in the format data-test-id=\"piano-key-’x’\", where x is the lowercase version of the corresponding keyboard character (e.g., the key for W uses data-test-id=\"piano-key-’w’\"). ### external_resources The corresponding audio files for each piano note are stored in the tunes directory, with filenames matching the lowercase key characters followed by .wav, such as tunes/a.wav, tunes/;.wav, tunes/d.wav, tunes/e.wav, and so on, covering all relevant keys including w, e, t, y, u, o, p, a, s, d, f, g, h, j, k, l, and ;. ### external_js_libraries No external JavaScript libraries are used in this application.   """fine_grained_reqs""" "1": "When the user clicks any visible piano key on the webpage, identified by its data-testid (e.g., piano-key-a for the white key a or piano-key-w for the black key w’), the system must play the corresponding sound file from the tunes directory (e.g., tunes/a.wav for key a’). The clicked key must receive the active class to visually highlight it, and the highlight must be removed after 150ms by removing the active class.", "3": ... "4": ... "5": ... "6": ...   """reference_answer""" https://github.com/SCUNLP/E2EDev/tree/main/E2ESD_Bench_06   """excutable_tests""" ### 1 # test_case Feature: Play piano sound and visually highlight the key when clicked - The system must play the corresponding piano sound and visually highlight the key when the user clicks on a piano key. The highlight should disappear after 150ms. Scenario: [Normal] User clicks on a white piano key - Given the webpage is loaded and the piano keys are visible - When the user clicks on the white piano key with data-testid "piano-key-a" - Then the system must play the sound - And the key with data-testid "piano-key-a" must have the "active" class added - And the "active" class must be removed from the key with data-testid "piano-key-a" after 150ms. # step_code from behave import given, when, then from selenium import webdriver from selenium.webdriver.common.by import By from selenium.webdriver.support.ui import WebDriverWait from selenium.webdriver.support import expected_conditions as EC import time @given(’the webpage is loaded and the piano keys are visible’) def step_given_webpage_loaded(context): context.driver = webdriver.Chrome() context.driver.get(f"file://{file_path}") context.driver.maximize_window() WebDriverWait(context.driver, 10).until( EC.visibility_of_element_located((By.CSS_SELECTOR, "[data-testid=’piano-key-a’]")) ) time.sleep(1) # Allow time for the page to fully load @when(’the user clicks on the white piano key with data-testid "piano-key-a"’) def step_when_user_clicks_piano_key(context): piano_key = WebDriverWait(context.driver, 10).until( EC.element_to_be_clickable((By.CSS_SELECTOR, "[data-testid=’piano-key-a’]")) ) piano_key.click() @then(’the system must play the sound’) def step_then_system_plays_sound(context): pass @then(’the key with data-testid "piano-key-a" must have the "active" class added’) def step_then_key_has_active_class(context): piano_key = context.driver.find_element(By.CSS_SELECTOR, "[data-testid=’piano-key-a’]") class_list = piano_key.get_attribute("class").split() assert "active" in class_list, "The active class was not added to the key" @then(’the "active" class must be removed from the key with data-testid "piano-key-a" after 150ms’) def step_then_active_class_removed(context): time.sleep(0.15) piano_key = context.driver.find_element(By.CSS_SELECTOR, "[data-testid=’piano-key-a’]") class_list = piano_key.get_attribute("class").split() assert "active" not in class_list, "The active class was not removed from the key" def after_scenario(context, scenario): if hasattr(context, driver’): context.driver.quit() # test_case ... # step_code ... # test_case ... # step_code ... ### 3 ... ### 4 ... ### 5 ... ### 6 ...
@@ B.1 Implementation of Baselines
LLM-based Frameworks for End-to-End Software Development. Recent advances in E2ESD using LLMs primarily fall into two categories: multi-agent and single-agent frameworks. The majority of existing work adopts a multi-agent paradigm, where the overall development process is decomposed into subtasks based on classical software engineering models (Sommerville, 2011). Within this category, Self-Collaboration (Dong et al., 2024) mimics the fundamental development stages—requirement analysis, implementation, and testing—via specialized agents. More structured approaches (Hong et al., 2024; Du et al., 2024; Rasheed et al., 2024; Sami et al., 2024) leverage the Waterfall model (Royce, 1987) to coordinate agent collaboration. Beyond this, MapCoder (Islam et al., 2024) incorporates prior project knowledge to improve software generation. In addition, while most frameworks rely on predefined workflows, recent efforts (Qian et al., 2024; Lin et al., 2024) explore dynamic agent interactions. ChatDev (Qian et al., 2024), for instance, enables agents to engage in multi-turn dialogues to iteratively refine the software product. In contrast, a smaller body of work pursues a single-agent approach, which assumes that a single LLM has sufficient capacity to independently complete the entire development cycle without explicit task decomposition (Engineer, 2024).
Implementations of LLM-based Frameworks. We reproduced six LLM-based code generation frameworks using their official GitHub implementations (if any): Vanilla LLM, GPT-Engineer, Self-Collaboration Code Generation, MapCoder, ChatDev, and MetaGPT.
  • For the Vanilla LLM baseline, we directly fed the task prompt (i.e., user requirements) together with the following structured instruction:
    Please wrap each part of the implementation in its appropriate Markdown code block:
    - Use """html for the HTML structure in **index.html**.
    - Use """css for the CSS styles in **style.css**.
    - Use """javascript for the JavaScript functionality in **script.js**.
    Ensure that the files are named exactly as indicated: **index.html**, **style.css**, and **script.js**.
    This will help maintain a consistent structure and make the code easily integrable.
    This instruction guides the model to generate modularized web projects containing HTML, CSS, and JavaScript components.
  • For GPT-Engineer and MetaGPT, we used their official implementations In particular, we employed their corresponding official Python packages, readily available for import and use. During our experiments, we executed the frameworks directly by feeding them the task prompts. No modifications were made to their internal workflows.
  • For the remaining frameworks—Self-Collaboration Code Generation, MapCoder, and ChatDev—we implemented all baselines using their publicly available source code from GitHub. However, both Self-Collaboration and MapCoder were originally designed for Python-only tasks. To adapt them to our multi-language (HTML/CSS/JavaScript) setting, we modified their agent system prompts accordingly.
    Specifically, for Self-Collaboration, we replaced the default PYTHON_DEVELOPER role with an H5_DEVELOPER role, enabling the agent to handle web development tasks:
@@
    A similar modification was applied to MapCoder, adapting its original Python-specific prompts to support full-stack web development. These adaptations ensured fair comparisons across all frameworks while preserving the core design principles of each system.

### v4

@@ 5 Conclusion
Dataset Scale and Construction Complexity. Unlike function-level tasks, repository-level coding tasks provide a more comprehensive assessment of LLMs’ capabilities in project construction and requirement fulfillment. However, constructing robust repository-level benchmarks remains challenging. A primary difficulty arises from the diverse and often idiosyncratic implementations that LLMs generate for identical user requirements, which substantially increases the cost and complexity of reliable automated evaluation. As a result, existing repository-level benchmarks, including E2EDev, are necessarily limited in scale (Table 1).
Scope Validity and Evaluation Protocol. Our strategic focus on Web applications is grounded in three methodological imperatives. First, this domain serves as a necessary "lower bound" for assessing E2ESD capabilities; the failures observed in this relatively deployable scenario reflect intrinsic limitations in LLM reasoning or framework design rather than domain-specific artifacts, suggesting that models struggling here are unlikely to succeed in more complex system-level settings. Second, to ensure rigorous reproducibility, we exclude CI/CD workflows and deep backend inspections, which rely heavily on volatile environmental dependencies that hinder consistent measurement. Third, aligned with Requirements Engineering standards (ISO/IEC/IEEE 29148), we adopt a black-box testing approach that verifies system correctness through user-perceptible behaviors via browser-based automation. This protocol allows us to validate whether the integrated system truly satisfies user intent without the noise of environment-specific internal implementation details.
Despite these limitations, our work serves as a reasonable starting point for more rigorous evaluation of E2ESD tasks. E2EDev contributes a standardized benchmark by providing explicit requirement specifications and reliable evaluation protocols, enabling systematic analysis of requirement fidelity and code consistency. While LLMs perform well on function-level and isolated programming tasks, our results reveal substantial gaps in scaling these capabilities to fully automated E2ESD settings. Given the significant resource investment required for benchmark construction, we plan to continuously expand E2EDev and maintain a public leaderboard to support reproducible and longitudinal evaluation.
Ethical Considerations
@@
Our work focuses solely on constructing the E2EDev benchmark for end-to-end requirement-to-executable code generation to evaluate the capabilities of LLM-based E2ESD frameworks. The study does not involve human subjects, sensitive personal data, or other ethical risks, and all source data are obtained from open-source communities. Any human annotation or verification of data was performed under controlled conditions. Human annotators were internal CS Master’s and PhD students, with priority given to those who had taken software testing or development courses. They were invited via email, participation was voluntary and unpaid, and written instructions explained the tasks and their intended use. Details are provided in Appendix A. All experiments and analyses adhere to the ACL Code of Ethics, and no procedures, datasets, or methods in this study raise concerns related to privacy, bias, or safety.
Reproducibility
@@
To facilitate reproducibility, we publicly release the full E2EDev benchmark source code on GitHub (https://github.com/SCUNLP/E2EDev), along with the complete benchmark dataset on Hugging Face (https://huggingface.co/datasets/GuanZhiZhao/E2EDev). Detailed setup and usage instructions are provided in both Appendix D.3 and the repository README. All experiments reported in this paper, including dataset processing, model evaluation, and metric computation, can be fully reproduced using these resources.
Artifacts License and Usage Terms
All code and data used and created in this work comply with their respective licenses and usage terms. External code and datasets are open-source or publicly available, and their licenses (e.g., CC-BY 4.0 for data, MIT for code) are respected, including proper attribution to the original creators. The E2EDev benchmark we create is released under CC-BY 4.0 for data and MIT for code, and is intended for research purposes.
LLMs Usage
Acknowledgments
@@
This work was supported in part by the National Natural Science Foundation of China (No. U25B201508, No. 62272330, and No.U24A20328); in part by the Singapore Ministry of Education (MOE) Academic Research Fund (AcRF) Tier 1 grant (Proposal ID: 24-SIS-SMU-002).
@@ A.1 Details on Human Verification for Project Selection
We select source projects through a two-step process. First, we crawl GitHub repositories using keywords such as “mini project”, “software project”, “web application”, and “H5”, retaining only those with more than 500 stars.
@@ A.2.1 HITL-MAA Annotation Process
A.2.2 Protocols for Human-in-the-Loop & Manual Verification
Manual Verification Protocol & Inner-Agreement. To ensure the consistency and quality of free-text requirements and test case validation, we implement a post-hoc gold evaluation process. Five human annotators with expertise in software testing participate, with one senior expert designated as the Gold Checker, responsible for reviewing others’ annotations. This process employs an iterative annotation and feedback mechanism to resolve inconsistencies efficiently. For both requirements and test cases, four regular annotators first perform validation. For requirement annotations, annotators interact with the web application through a browser to directly experience and verify functionality. For test cases, written in natural language as step-by-step operational guides, annotators follow the instructions to ensure consistency and validity against the requirements. Detected discrepancies or errors are revised before submission to the Gold Checker. The Gold Checker evaluates each annotation on three binary criteria: equivalence, completeness, and correctness, where a score of 1 indicates the criterion is met and 0 otherwise. Only when the Fleiss’ Kappa score for all three aspects exceeds 0.7 does the Gold Checker consolidate a final version of the requirement or test case, which may include discarding it if necessary. If the agreement falls below the threshold, the Gold Checker provides detailed feedback, and annotators repeat the annotation process. Newly proposed test cases also undergo gold evaluation. In such cases, all three criteria are initially scored as 0 by the annotators, and the Gold Checker reviews the new case before returning it to all regular annotators for reconsideration and re-annotation. If annotators fail to recognize a new test case, the Gold Checker similarly returns it for re-annotation. This approach balances thoroughness and efficiency, resulting in a reliable, scalable, and reproducible manual annotation protocol suitable for complex free-text requirement and test case validation tasks. It ensures high-quality, consistent annotations for downstream applications.
Reliability Audit & Scalability Verification. To further mitigate potential bias from the initial team size and validate the generalizability of our quality standards, we conducted an extended reliability audit. We expanded the expert pool to 10 evaluators (adding five distinct external experts to the original group) and performed a blind review on a stratified random sample of 100 test cases from the final dataset. This broader panel independently assessed the sample under the same strict criteria. The resulting inter-rater agreement yielded a Fleiss’ Kappa of 0.81, demonstrating that our original 5-expert protocol aligns with the consensus of a larger expert body and ensuring the dataset’s high quality is statistically robust.
Input: Source code annotated with test-ids
Output: Executable and validated end-to-end test scripts with aligned requirements
1 /* Code Analyzer */;
@@ 6   Analyze logic functionality and connections with frontend elements;
8Summarize the analysis for downstream modules;
9
10/* Requirement Extractor */;
@@ 11 Extract contextual requirements using both source code and analysis summary;
13
14foreach req in user_req_lst do
15   Human annotator validates and corrects the extracted requirement;
@@ A.2.3 Analysis on Human Labor during HITL-MAA
Table 5: Comparison of Manual vs. Semi-Automated Annotation
@@ A.3 Case Studies of E2EDev
E2ESD_Bench_06 """prompt""" You are tasked with implementing a complete web application using HTML, JavaScript, and CSS. Your implementation must strictly follow the specifications described below. ### SUMMARY # overview The Playable Piano web application allows users to interact with a virtual piano. Users can click on piano keys or press corresponding keyboard keys to play notes. The application includes features for adjusting volume and toggling the visibility of key labels. # predefined_options The web application initializes with a default volume level set to 0.5, and piano key labels are displayed by default upon launch. It features a virtual piano with both black and white keys, each mapped to specific keyboard characters. The black keys correspond to the uppercase letters W, E, T, Y, U, O, and P, while the white keys correspond to A, S, D, F, G, H, J, K, L, and ;. Each key element in the DOM is assigned a unique data-test-id attribute in the format data-test-id=\"piano-key-’x’\", where x is the lowercase version of the corresponding keyboard character (e.g., the key for W uses data-test-id=\"piano-key-’w’\"). # external_resources The corresponding audio files for each piano note are stored in the tunes directory, with filenames matching the lowercase key characters followed by .wav, such as tunes/a.wav, tunes/;.wav, tunes/d.wav, tunes/e.wav, and so on, covering all relevant keys including w, e, t, y, u, o, p, a, s, d, f, g, h, j, k, l, and ;. # external_js_libraries No external JavaScript libraries are used in this application. ### Functional Requirements Implement the following features as described. For each requirement, make sure the HTML structure, JavaScript behavior, and CSS styles match the specifications exactly. # REQUIREMENTS: - Requirement 1: When the user clicks any visible piano key on the webpage, identified by its data-testid (e.g., piano-key-a for the white key a or piano-key-w for the black key w’), the system must play the corresponding sound file from the tunes directory (e.g., tunes/a.wav for key a’). The clicked key must receive the active class to visually highlight it, and the highlight must be removed after 150ms by removing the active class. - Requirement 2:... - Requirement 3:... - Requirement 4:... - Requirement 5:...   """requirement_summary""" ### overview The Playable Piano web application allows users to interact with a virtual piano. Users can click on piano keys or press corresponding keyboard keys to play notes. The application includes features for adjusting volume and toggling the visibility of key labels. ### predefined_options The web application initializes with a default volume level set to 0.5, and piano key labels are displayed by default upon launch. It features a virtual piano with both black and white keys, each mapped to specific keyboard characters. The black keys correspond to the uppercase letters W, E, T, Y, U, O, and P, while the white keys correspond to A, S, D, F, G, H, J, K, L, and ;. Each key element in the DOM is assigned a unique data-test-id attribute in the format data-test-id=\"piano-key-’x’\", where x is the lowercase version of the corresponding keyboard character (e.g., the key for W uses data-test-id=\"piano-key-’w’\"). ### external_resources The corresponding audio files for each piano note are stored in the tunes directory, with filenames matching the lowercase key characters followed by .wav, such as tunes/a.wav, tunes/;.wav, tunes/d.wav, tunes/e.wav, and so on, covering all relevant keys including w, e, t, y, u, o, p, a, s, d, f, g, h, j, k, l, and ;. ### external_js_libraries No external JavaScript libraries are used in this application.   """fine_grained_reqs""" "1": "When the user clicks any visible piano key on the webpage, identified by its data-testid (e.g., piano-key-a for the white key a or piano-key-w for the black key w’), the system must play the corresponding sound file from the tunes directory (e.g., tunes/a.wav for key a’). The clicked key must receive the active class to visually highlight it, and the highlight must be removed after 150ms by removing the active class.", "3": ... "4": ... "5": ... "6": ...   """reference_answer""" https://github.com/GuanZhizhao/E2EDev/tree/main/E2ESD_Bench_06   """excutable_tests""" ### 1 # test_case Feature: Play piano sound and visually highlight the key when clicked - The system must play the corresponding piano sound and visually highlight the key when the user clicks on a piano key. The highlight should disappear after 150ms. Scenario: [Normal] User clicks on a white piano key - Given the webpage is loaded and the piano keys are visible - When the user clicks on the white piano key with data-testid "piano-key-a" - Then the system must play the sound - And the key with data-testid "piano-key-a" must have the "active" class added - And the "active" class must be removed from the key with data-testid "piano-key-a" after 150ms. # step_code from behave import given, when, then from selenium import webdriver from selenium.webdriver.common.by import By from selenium.webdriver.support.ui import WebDriverWait from selenium.webdriver.support import expected_conditions as EC import time @given(’the webpage is loaded and the piano keys are visible’) def step_given_webpage_loaded(context): context.driver = webdriver.Chrome() context.driver.get(f"file://{file_path}") context.driver.maximize_window() WebDriverWait(context.driver, 10).until( EC.visibility_of_element_located((By.CSS_SELECTOR, "[data-testid=’piano-key-a’]")) ) time.sleep(1) # Allow time for the page to fully load @when(’the user clicks on the white piano key with data-testid "piano-key-a"’) def step_when_user_clicks_piano_key(context): piano_key = WebDriverWait(context.driver, 10).until( EC.element_to_be_clickable((By.CSS_SELECTOR, "[data-testid=’piano-key-a’]")) ) piano_key.click() @then(’the system must play the sound’) def step_then_system_plays_sound(context): pass @then(’the key with data-testid "piano-key-a" must have the "active" class added’) def step_then_key_has_active_class(context): piano_key = context.driver.find_element(By.CSS_SELECTOR, "[data-testid=’piano-key-a’]") class_list = piano_key.get_attribute("class").split() assert "active" in class_list, "The active class was not added to the key" @then(’the "active" class must be removed from the key with data-testid "piano-key-a" after 150ms’) def step_then_active_class_removed(context): time.sleep(0.15) piano_key = context.driver.find_element(By.CSS_SELECTOR, "[data-testid=’piano-key-a’]") class_list = piano_key.get_attribute("class").split() assert "active" not in class_list, "The active class was not removed from the key" def after_scenario(context, scenario): if hasattr(context, driver’): context.driver.quit() # test_case ... # step_code ... # test_case ... # step_code ... ### 3 ... ### 4 ... ### 5 ... ### 6 ...
@@ B.1 Implementation of Baselines
We select a diverse set of representative requirement-to-executable software development frameworks across various LLM backbones. These include: (1) Vanilla LLM that generates the full codebase via simply prompting LLM using user requirements without external scaffolding. (2) Single-Agent Framework, GPT-Engineer (Engineer, 2024), which builds the complete codebase in a single reasoning pass with minimal external coordination while maintaining a modular development workflow. (3) Multi-Agent Frameworks, a dominant paradigm, where the development process is decomposed into specialized agent roles based on classic software engineering principles (Sommerville, 2011). We benchmark with a wide range of state-of-the-art multi-agent framework, including Self-Collaboration (Dong et al., 2024), MetaGPT (Hong et al., 2024), MapCoder (Islam et al., 2024), and ChatDev (Qian et al., 2024). We additionally include two widely used coding agents, OpenHands (Wang et al., 2024) and SWE-Agent (Yang et al., 2024), as reference systems. These agents are designed for general-purpose software engineering scenarios with extensive tool and environment interaction, which differ from the requirement-to-executable software generation setting considered in E2EDev. Their results are reported in the Appendix  C.1.
Finally, we adopt a range of backbones from two vendors (Hurst et al., 2024; Hui et al., 2024), covering various scales and architectures: Claude-Haiku 4.5, GPT-4o, GPT-4o-mini, Qwen-7B (Qwen2.5-7B-Instruct), Qwen-70B (Qwen2.5-72B-Instruct), and the mixture-of-experts Qwen-Max (Qwen2.5-Max). This enables controlled analysis of framework versus model backbone effects. Implementation details are presented below.
Implementations of LLM-based Frameworks. We reproduced eight LLM-based code generation frameworks using their official GitHub implementations (if available): Vanilla LLM, GPT-Engineer, Self-Collaboration Code Generation, MapCoder, ChatDev, MetaGPT, OpenHands, and SWE-Agent.
Vanilla LLM. We directly fed the task prompts (i.e., user requirements) along with the following structured instruction:
@@
GPT-Engineer. GPT-Engineer builds a complete codebase in a single reasoning pass with minimal external coordination, maintaining a modular workflow. We used the official Python packages and executed the frameworks directly with task prompts, without modifying internal workflows.
Self-Collaboration Code Generation. This framework incorporates software-development methodology by assembling a team of three LLM roles—analyst, coder, and tester—responsible for analysis, coding, and testing. Originally designed for Python-only tasks, we adapted it to our multi-language (HTML/CSS/JavaScript) setting by modifying the system prompts. Specifically, we replaced the default PYTHON_DEVELOPER role with an H5_DEVELOPER role:
H5_DEVELOPER = """
I want you to act as an H5 developer on our development team. ...,
write HTML, JavaScript, and CSS code ... provide the updated HTML,
JavaScript, and CSS code ...
MapCoder. MapCoder consists of four agents simulating the stages of software development: recalling examples, planning, code generation, and debugging. Similar to Self-Collaboration, it was originally Python-focused; we adapted it to support HTML/CSS/JavaScript by modifying its agent system prompts accordingly.
ChatDev. ChatDev is a chat-powered software development framework where LLM-driven agents communicate via structured chat chains. Agents contribute to design, coding, and testing phases through multi-turn dialogues, guided by communicative strategies to reduce hallucinations. We implemented it using the official GitHub source without modifications.
MetaGPT. MetaGPT is a meta-programming framework that incorporates human workflow principles into LLM-based multi-agent collaboration. It encodes Standardized Operating Procedures (SOPs) into prompt sequences, enabling agents with domain expertise to verify intermediate results and reduce errors. Using an assembly-line paradigm, MetaGPT assigns diverse roles to multiple agents to efficiently decompose complex tasks. We used the official implementations and Python packages, executing them directly with task prompts without internal modifications.
OpenHands. OpenHands is an interactive human–agent software development framework that supports multi-turn collaboration for iterative project construction. The system relies on continuous human feedback and intervention throughout the development process, which makes large-scale automated evaluation time-consuming. To align it as closely as possible with the E2EDev setting, we restrict OpenHands to a single execution without any human interaction. We use the official GitHub implementation without modification and report the results in Appendix C.1.
SWE-Agent. SWE-Agent is designed for repository-level software engineering tasks, where the goal is to resolve GitHub issues by modifying an existing codebase, as evaluated in the SWE-bench setting. Its input consists of an existing repository and a natural language issue description, and the output is a patched repository rather than a newly generated project. For our setting, we replace the original prompt template—which instructs the model to read a project from {{repo_path}} and modify it according to {{issue}}—with a new template that directs the model to generate a runnable project directly from the given user requirement. We implement SWE-Agent using the official GitHub source without further modification, and report its results in Appendix C.1.
Evaluation Coverage. We consider eight LLM-based requirement-to-executable software development frameworks in total. Six frameworks (Vanilla LLM, GPT-Engineer, Self-Collaboration, MapCoder, ChatDev, and MetaGPT) are evaluated across all six LLM backbones. Due to their different design goals and substantially higher evaluation cost, OpenHands and SWE-Agent are evaluated only on Claude-Haiku-4.5 and GPT-4o, with their results reported in Appendix C.1.